# The algebraic and geometric classification of

### v1

@@ 1.3.10. Compatible pre-Lie algebras defined on $\mathcal{C}_{06}^{0}$
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2},$&$e_{1}\ast e_{2}=\alpha_{3}e_{2},$&$e_{2}\ast e_{1}=\alpha_{1}e_{2}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{1}=e_{2},$&&\\ $e_{1}\ast e_{1}=\alpha_{4}e_{1},$&$e_{1}\ast e_{2}=\alpha_{5}e_{1},$&$e_{2}\ast e_{1}=\alpha_{4}e_{2},$&$e_{2}\ast e_{2}=\alpha_{5}e_{2}.$\end{tabular}\right.$
We may assume $\alpha_{5}\neq 0$.
Let us first consider the first form. Then we have the following cases:
@@
  • If $\alpha_{3}=0$, then we consider the following two cases.
@@
    • If $\alpha_{2}=0$, then we obtain the parametric family $\mathcal{C}_{32}^{0,\beta,0}$
Now, we consider the second form. Then choose $\xi={\alpha^{-1}_{5}},\nu=-{\alpha_{4}}{\alpha^{-1}_{5}}$ and obtain the algebra:
${\mathcal{C}_{33}}:=\left\{\begin{tabular}[]{ll}$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{2}=e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
1.3.11. Compatible pre-Lie algebras defined on $\mathcal{C}_{06}^{1}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{06}^{1},\mathcal{C}_{06}^{1})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{1}$ are of the forms:
@@ 1.3.11. Compatible pre-Lie algebras defined on $\mathcal{C}_{06}^{1}$
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2},$&$e_{1}\ast e_{2}=\alpha_{3}e_{2},$&$e_{2}\ast e_{1}=\alpha_{1}e_{2}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha_{4}e_{1},$&$e_{1}\ast e_{2}=\alpha_{4}e_{2},$&$e_{2}\ast e_{1}=\alpha_{4}e_{2},$&$e_{2}\ast e_{2}=\alpha_{5}e_{1}+\alpha_{6}e_{2}.$\end{tabular}\right.$
We may assume $\left(\alpha_{5},\alpha_{6}\right)\neq\left(0,0\right)$.
If the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{1}$ is of the first form, then we obtain the algebras $\mathcal{C}_{31}^{1,\beta,\gamma}$ and $\mathcal{C}_{32}^{1,\beta,\gamma}$.
@@
  • If $\alpha_{5}\neq 0$, then choose $\xi={\alpha^{-\frac{1}{2}}_{5}}$ and obtain the parametric family:
    $\mathcal{C}_{34}^{\alpha,\beta}:=\left\{\begin{tabular}[]{llll}$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=\alpha e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2},$&$e_{2}\ast e_{2}=e_{1}+\beta e_{2}.$\end{tabular}\right.$
@@
  • If $\alpha_{5}=0$, then choose $\xi={\alpha^{-1}_{6}}$ and obtain the parametric family:
    $\mathcal{C}_{35}^{\alpha}:=\left\{\begin{tabular}[]{llll}$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=\alpha e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
@@
1.3.12. Compatible pre-Lie algebras defined on $\mathcal{C}_{06}^{2}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{06}^{2},\mathcal{C}_{06}^{2})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{2}$ are of the forms:
@@ 1.3.12. Compatible pre-Lie algebras defined on $\mathcal{C}_{06}^{2}$
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=2e_{2},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2},$&$e_{1}\ast e_{2}=\alpha_{3}e_{2},$&$e_{2}\ast e_{1}=\alpha_{1}e_{2}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=2e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha_{4}e_{1}+\alpha_{5}e_{2},$&$e_{1}\ast e_{2}=2\alpha_{4}e_{2},$&$e_{2}\ast e_{1}=\alpha_{6}e_{1}+\alpha_{4}e_{2},$&$e_{2}\ast e_{2}=2\alpha_{6}e_{2}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=2e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha_{7}e_{1},$&$e_{1}\ast e_{2}=\alpha_{8}e_{1}+2\alpha_{7}e_{2},$&$e_{2}\ast e_{1}=2\alpha_{8}e_{1}+\alpha_{7}e_{2},$&$e_{2}\ast e_{2}=\alpha_{8}e_{2}.$\end{tabular}\right.$
We may assume $\alpha_{6}\alpha_{8}\neq 0$.
If the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{2}$ is of the first form, then we obtain the algebras $\mathcal{C}_{31}^{2,\beta,\gamma}$ and $\mathcal{C}_{32}^{2,\beta,\gamma}$.
Assume now that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{2}$ are of the second form. Then, choose $\xi={\alpha^{-1}_{6}}$ and obtain the parametric family:
$\mathcal{C}_{36}^{\alpha,\beta}:=\left\{\begin{tabular}[]{llll}$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=2e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{1}\ast e_{2}=2\alpha e_{2},$&$e_{2}\ast e_{1}=e_{1}+\alpha e_{2},$&$e_{2}\ast e_{2}=2e_{2}.$\end{tabular}\right.$
@@
Finally, assume that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{2}$ are of the third form. Then, choose $\xi={\alpha^{-1}_{8}}$ and obtain the parametric family:
$\mathcal{C}_{37}^{\alpha}:=\left\{\begin{tabular}[]{llll}$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=2e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=e_{1}+2\alpha e_{2},$&$e_{2}\ast e_{1}=2e_{1}+\alpha e_{2},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
@@
1.3.13. Compatible pre-Lie algebras defined on $\mathcal{C}_{07}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{07},\mathcal{C}_{07})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{07}$ are of the forms:
@@ 1.3.13. Compatible pre-Lie algebras defined on $\mathcal{C}_{07}$
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{2}=e_{2},$&&\\ $e_{1}\ast e_{1}=\left(\alpha_{3}+\alpha_{2}-\alpha_{1}\right)e_{1}-\alpha_{2}e_{2},$&$e_{1}\ast e_{2}=\alpha_{1}e_{1}+\alpha_{2}e_{2},$\\ $e_{2}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2},$&$e_{2}\ast e_{2}=-\alpha_{1}e_{1}+\alpha_{3}e_{2}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha_{4}e_{1},$&$e_{2}\ast e_{2}=\alpha_{5}e_{2}.$\end{tabular}\right.$
We may assume that $\alpha_{4}\neq\alpha_{5}$.
Suppose first that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{07}$ has the first form. Then we obtain the algebras:
$\mathcal{C}_{38}^{\alpha,\beta,\gamma}:=\left\{\begin{tabular}[]{llll}$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{2}=e_{2},$&&\\ $e_{1}\ast e_{1}=\left(\gamma+\beta-\alpha\right)e_{1}-\beta e_{2},$&$e_{1}\ast e_{2}=\alpha e_{1}+\beta e_{2},$&\\ $e_{2}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{2}\ast e_{2}=-\alpha e_{1}+\gamma e_{2}.$\end{tabular}\right.\qquad$
The algebras $\mathcal{C}_{38}^{\alpha,\beta,\gamma}$ and $\mathcal{C}_{38}^{\alpha^{\prime},\beta^{\prime},\gamma^{\prime}}$ are isomorphic if and only if
$(\alpha,\beta,\gamma)=(\alpha^{\prime},\beta^{\prime},\gamma^{\prime})$ or $(\alpha,\beta,\gamma)=(\beta^{\prime},\alpha^{\prime},-\alpha^{\prime}+\beta^{\prime}+\gamma^{\prime})$.
Now, assume that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{07}$ has the second form. Then we obtain the algebras:
$\mathcal{C}_{39}^{\alpha,\beta\neq\alpha}:=\left\{\begin{tabular}[]{ll}$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{2}\ast e_{2}=\beta e_{2}.$\end{tabular}\right.\qquad$
The algebras $\mathcal{C}_{39}^{\alpha,\beta}$ and $\mathcal{C}_{39}^{\alpha^{\prime},\beta^{\prime}}$ are isomorphic if and only if $(\alpha,\beta)=(\alpha^{\prime},\beta^{\prime})$ or $(\alpha,\beta)=(\beta^{\prime},\alpha^{\prime}).$
1.3.14. Compatible pre-Lie algebras defined on $\mathcal{C}_{08}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{08},\mathcal{C}_{08})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{08}$ are of the forms:

### v2

@@ 2.3.10 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{06}^{0}}$
    $e_{1}\cdot e_{1}=e_{1}$, $e_{2}\cdot e_{1}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{1}\ast e_{2}=\alpha_{3}e_{2}$, $e_{2}\ast e_{1}=\alpha_{1}e_{2}$,
@@
    $e_{1}\cdot e_{1}=e_{1}$, $e_{2}\cdot e_{1}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{4}e_{1}$, $e_{1}\ast e_{2}=\alpha_{5}e_{1}$, $e_{2}\ast e_{1}=\alpha_{4}e_{2}$, $e_{2}\ast e_{2}=\alpha_{5}e_{2}$.
We may assume $\alpha_{5}\neq 0$. Let us first consider the first form. Then we have the following cases:
@@
  • If $\alpha_{3}=0$, then we consider the following two cases:
@@
    • If $\alpha_{2}=0$, then we obtain the parametric family $\mathcal{C}_{32}^{0,\beta,0}$.
Now, we consider the second form. Then choose $\xi={\alpha^{-1}_{5}}$, $\nu=-{\alpha_{4}}{\alpha^{-1}_{5}}$ and obtain the algebra
${\mathcal{C}_{33}}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{2}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{2}=e_{1},\quad e_{2}\ast e_{2}=e_{2}.$
2.3.11 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{06}^{1}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}\bigl{(}\mathcal{C}_{06}^{1},\mathcal{C}_{06}^{1}\bigr{)}$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{1}$ are of the following forms:
@@ 2.3.11 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{06}^{1}}$
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=e_{2}$, $e_{2}\cdot e_{1}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{1}\ast e_{2}=\alpha_{3}e_{2}$, $e_{2}\ast e_{1}=\alpha_{1}e_{2}$,
@@
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=e_{2}$, $e_{2}\cdot e_{1}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{4}e_{1}$, $e_{1}\ast e_{2}=\alpha_{4}e_{2}$, $e_{2}\ast e_{1}=\alpha_{4}e_{2}$, $e_{2}\ast e_{2}=\alpha_{5}e_{1}+\alpha_{6}e_{2}$.
We may assume $(\alpha_{5},\alpha_{6})\neq(0,0)$. If the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{1}$ is of the first form, then we obtain the algebras $\mathcal{C}_{31}^{1,\beta,\gamma}$ and $\mathcal{C}_{32}^{1,\beta,\gamma}$.
@@
  • If $\alpha_{5}\neq 0$, then choose $\xi={\alpha^{-\frac{1}{2}}_{5}}$ and obtain the parametric family
    $\mathcal{C}_{34}^{\alpha,\beta}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
    $\ e_{1}\ast e_{1}=\alpha e_{1},\quad e_{1}\ast e_{2}=\alpha e_{2},\quad e_{2}\ast e_{1}=\alpha e_{2},\quad e_{2}\ast e_{2}=e_{1}+\beta e_{2}.$
@@
  • If $\alpha_{5}=0$, then choose $\xi={\alpha^{-1}_{6}}$ and obtain the parametric family
    $\mathcal{C}_{35}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
    $\ e_{1}\ast e_{1}=\alpha e_{1},\quad e_{1}\ast e_{2}=\alpha e_{2},\quad e_{2}\ast e_{1}=\alpha e_{2},\quad e_{2}\ast e_{2}=e_{2}.$
@@
2.3.12 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{06}^{2}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}\bigl{(}\mathcal{C}_{06}^{2},\mathcal{C}_{06}^{2}\bigr{)}$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{2}$ are of the following forms:
@@ 2.3.12 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{06}^{2}}$
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=2e_{2}$, $e_{2}\cdot e_{1}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{1}\ast e_{2}=\alpha_{3}e_{2}$, $e_{2}\ast e_{1}=\alpha_{1}e_{2}$,
@@
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=2e_{2}$, $e_{2}\cdot e_{1}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{4}e_{1}+\alpha_{5}e_{2}$, $e_{1}\ast e_{2}=2\alpha_{4}e_{2}$, $e_{2}\ast e_{1}=\alpha_{6}e_{1}+\alpha_{4}e_{2}$, $e_{2}\ast e_{2}=2\alpha_{6}e_{2}$,
@@
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=2e_{2}$, $e_{2}\cdot e_{1}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{7}e_{1}$, $e_{1}\ast e_{2}=\alpha_{8}e_{1}+2\alpha_{7}e_{2}$, $e_{2}\ast e_{1}=2\alpha_{8}e_{1}+\alpha_{7}e_{2}$, $e_{2}\ast e_{2}=\alpha_{8}e_{2}$.
We may assume $\alpha_{6}\alpha_{8}\neq 0$. If the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{2}$ is of the first form, then we obtain the algebras $\mathcal{C}_{31}^{2,\beta,\gamma}$ and $\mathcal{C}_{32}^{2,\beta,\gamma}$.
Assume now that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{2}$ are of the second form. Then, choose $\xi={\alpha^{-1}_{6}}$ and obtain the parametric family
$\mathcal{C}_{36}^{\alpha,\beta}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=2e_{2},\quad e_{2}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},$
$\ e_{1}\ast e_{2}=2\alpha e_{2},\quad e_{2}\ast e_{1}=e_{1}+\alpha e_{2},\quad e_{2}\ast e_{2}=2e_{2}.$
@@
Finally, assume that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{2}$ are of the third form. Then, choose $\xi={\alpha^{-1}_{8}}$ and obtain the parametric family
$\mathcal{C}_{37}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=2e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
$\ e_{1}\ast e_{1}=\alpha e_{1},\quad e_{1}\ast e_{2}=e_{1}+2\alpha e_{2},\quad e_{2}\ast e_{1}=2e_{1}+\alpha e_{2},\quad e_{2}\ast e_{2}=e_{2}.$
@@
2.3.13 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{07}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{07},\mathcal{C}_{07})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{07}$ are of the following forms:
@@ 2.3.13 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{07}}$
    $e_{1}\cdot e_{1}=e_{1}$, $e_{2}\cdot e_{2}=e_{2}$, $e_{1}\ast e_{1}=(\alpha_{3}+\alpha_{2}-\alpha_{1})e_{1}-\alpha_{2}e_{2}$, $e_{1}\ast e_{2}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{2}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{2}\ast e_{2}=-\alpha_{1}e_{1}+\alpha_{3}e_{2}$,
@@
    $e_{1}\cdot e_{1}=e_{1}$, $e_{2}\cdot e_{2}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{4}e_{1}$, $e_{2}\ast e_{2}=\alpha_{5}e_{2}$.
We may assume that $\alpha_{4}\neq\alpha_{5}$. Suppose first that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{07}$ has the first form. Then we obtain the algebras
$\mathcal{C}_{38}^{\alpha,\beta,\gamma}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{2}\cdot e_{2}=e_{2},\quad e_{1}\ast e_{1}=(\gamma+\beta-\alpha)e_{1}-\beta e_{2},$
$\ e_{1}\ast e_{2}=\alpha e_{1}+\beta e_{2},\quad e_{2}\ast e_{1}=\alpha e_{1}+\beta e_{2},\quad e_{2}\ast e_{2}=-\alpha e_{1}+\gamma e_{2}.$
The algebras $\mathcal{C}_{38}^{\alpha,\beta,\gamma}$ and $\mathcal{C}_{38}^{\alpha^{\prime},\beta^{\prime},\gamma^{\prime}}$ are isomorphic if and only if $(\alpha,\beta,\gamma)=\bigl{(}\alpha^{\prime},\beta^{\prime},\gamma^{\prime}\bigr{)}$ or $(\alpha,\beta,\gamma)=\bigl{(}\beta^{\prime},\alpha^{\prime},-\alpha^{\prime}+\beta^{\prime}+\gamma^{\prime}\bigr{)}$.
Now, assume that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{07}$ has the second form. Then we obtain the algebras
$\mathcal{C}_{39}^{\alpha,\beta\neq\alpha}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{2}\cdot e_{2}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{1},\quad e_{2}\ast e_{2}=\beta e_{2}.$
The algebras $\mathcal{C}_{39}^{\alpha,\beta}$ and $\mathcal{C}_{39}^{\alpha^{\prime},\beta^{\prime}}$ are isomorphic if and only if $(\alpha,\beta)=\bigl{(}\alpha^{\prime},\beta^{\prime}\bigr{)}$ or $(\alpha,\beta)=\bigl{(}\beta^{\prime},\alpha^{\prime}\bigr{)}$.
2.3.14 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{08}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{08},\mathcal{C}_{08})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{08}$ are of the following forms: